A 50kg track athlete running at 10.0 m/s jumps over a hurdle. If athlete’s horizontal component of the velocity over the hurdle is 8.8 m/s and the air resistance is disregarded, how high was the jump?
1
Expert's answer
2020-04-08T10:30:00-0400

As per the given question,

Mass of the athlete (m)=50 kg

Speed of the athlete (v)=10 m/sec

Let the angle of projection "=\theta"

Let the jump was h meter high.

Horizontal component of the velocity "=v \cos\theta=8.8 m/sec"

"\Rightarrow 10 \cos\theta =8.8"

"\Rightarrow \cos \theta =\dfrac{8.8}{10}=0.88"

"\Rightarrow \theta =28.35^\circ"

Vertical component of the velocity"=v\sin\theta=10 \sin 28.35^\circ"

"\Rightarrow h=\dfrac{v^2\sin^2\theta}{2g}"

"\Rightarrow h=\dfrac{10\times 10 \sin^2 28.35^\circ}{9.8}"

"\Rightarrow h=2.3 m"